OpenStudy (anonymous):

What is the standard form of ( 5i ) (6-5i ) ---------- ? 2i(-3+7i) A) -111+126i --------- 40 B) -94+119i --------- 40 C)-94 +119i ------------ 58 D) 94-119i -------- 58

OpenStudy (anonymous):

start by multiplying out in your numerator and denominator to get: \[\frac{ (5i)(6-5i) }{ (2i)(-3+7i) } = \frac{ 30i-25i^2 }{ -6i+17i^2 } = \frac{25+ 30i}{-17-6i}\] then we need to simplify the bottom term. We can do this by multiplying by its complex conjugate \[\frac{25+ 30i}{-17-6i}\ * \frac{ -17+6i }{ -17+6i }\] try and finish from here!